A block of glass with flat sides is surrounded by air. A ray of light traveling in the glass (refractive index na) is incident on the flat interface with air at the top surface of the block of glass. You observe that no light is refracted into the air if the incident ray in the glass has an angle of incidence (measured relative to the normal to the surface of the glass) greater than 35.5°. What is the refractive index of the glass?
Answer #1

Here ,

crictical angle , theta = 35.5 degree

let the refractive index is na

Using snell's law

na * sin(thetaC) = n1 * sin(90)

na * sin(35.5) = 1 * sin(90)

na = 1/sin(35.5)

na = 1.722

the refractive index of the glass is 1.722